I need help with E, F, and G. I have provided the answers for parts A-D.

Zan Azlett and Angela Zesiger have joined forces to start​ A&Z Lettuce​ Products, a processor of packaged shredded lettuce for institutional use. Zan has years of food processing​ experience, and Angela has extensive commercial food preparation experience. The process will consist of opening crates of lettuce and then​ sorting, washing,​ slicing, preserving, and finally packaging the prepared lettuce.​ Together, with help from​ vendors, they think they can adequately estimate​ demand, fixed​ costs, revenues, and variable cost per bag of lettuce. They think a largely manual process will have monthly fixed costs of $36,000 and variable costs of $2.00 per bag. A more mechanized process will have fixed costs of $72,000 per month with variable costs of $1.00 per bag. They expect to sell the shredded lettuce for $2.50 per bag.

​a) The​ break-even quantity in units for the manual process​ =

72000

bags ​(round your response to the nearest whole​ number).

​b) The revenue for the manual process at the​ break-even quantity​ =

$180000

​(round your response to the nearest whole​ number).

​c) The​ break-even quantity in units for the mechanized process​ =

$48000

bags ​(round your response to the nearest whole​ number).

​d) The revenue for the mechanized process at the​ break-even quantity​ =

120000

​(round your response to the nearest whole​ number).

​e) For monthly sales of 55,000 ​bags, for the option with manua l​processing, A&Z Lettuce Products will have a profit of​(round your response to the nearest whole number and include a minus sign if the profit is​negative).

​f) For monthly sales of 55,000 ​bags, for the option with mechanized ​processing, A&Z Lettuce Products will have a profit of​​(round your response to the nearest whole number and include a minus sign if the profit is​ negative).

​g) The quantity at which Zan and Angela are going to be indifferent between the manual and mechanized process​ = bags​(round your response to the nearest whole​ number).
Answer #1

In manual process, contribution per unit is 2.5-2 =0.5 per bag.

Profit with 55000 bags = 55000*0.5-36000= -8500

In mechanised process, contribution = 2.5-1 = 1.5 per bag

Profit with the process = 1.5*55000-72000 = 10500

G.

If n is the number at which cost is equal for both options

1.5*n+72000= 0.5*n+36000

n= 36000 units